A bond has a quoted price of $900 today. It has a face value of $1,000, coupon rate of 12%, and a time to maturity of 6 years. Coupons are paid every month. What is its yield to maturity?
Answer #1

Coupon = (12% of 1000) / 12 = 10

Number of periods = 12 * 6 = 72

Yield to maturity = 14.51%

Keys to use in a financial calculator:

2nd P/Y 12

FV 1000

PV -900

PMT 10

N 72

CPT I/Y